40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  ATF eLog  Not logged in ELOG logo
Entry  Mon Feb 7 11:57:38 2022, Stephen, DailyProgress, Cryo vacuum chamber, Doubled Thermal Linkage Capacity run started 11 Feb 2022 D2000310_y-004_section_double_linkage_20220207.png
    Reply  Tue Feb 15 11:59:22 2022, Radhika, DailyProgress, Cryo vacuum chamber, Doubled Thermal Linkage Capacity run started 11 Feb 2022 Cu_bar_comparison.pdfcooldown_0211_all.pdf
       Reply  Tue Feb 15 16:54:27 2022, rana, DailyProgress, Cryo vacuum chamber, Doubled Thermal Linkage Capacity run started 11 Feb 2022 
    Reply  Wed Feb 16 11:35:18 2022, Stephen, DailyProgress, Cryo vacuum chamber, Doubled Thermal Linkage Capacity run started 11 Feb 2022 
Message ID: 2717     Entry time: Mon Feb 7 11:57:38 2022     Reply to this: 2720   2722
Author: Stephen 
Type: DailyProgress 
Category: Cryo vacuum chamber 
Subject: Doubled Thermal Linkage Capacity run started 11 Feb 2022 

[Stephen, Radhika]

The double copper bar configuration pictured in Attachment 1 has been implemented. We completed the updates within work sessions on Thursday and Friday. Here's a pseudo-log:

  • Removed all items from coldplate to make space for thermal linkage.
  • Prepared thermal linkage for installation.
    • ref. QIL/2694 for all preparations made.
  • Installed thermal linkage.
    • Removed thermal linkage to allow for regreasing
      • Discovered a single brass #4 washer sandwiched between coldhead and rigid copper bar (likely lost during efforts to install an RTD during intitial thermal linkage install); thermal grease was mostly indented at brass washer and at area opposite location of brass washer, suggesting non-uniform contact.
    • Applied Apiezon N thermal grease at joints: [coldhead - top copper bar], [top copper bar - bottom copper bar, near coldhead], [top copper bar - bottom copper bar, near thermal strap].
    • Sandwiched together top and bottom copper bars, to insert simultaneously.
    • Angled top and bottom copper bars to pass below coldhead, and overall passing the two bars in together went smoothly.
    • Bolted both copper bars into coldhead using long (1.25") 4-40 bolts threaded into coldhead. Access was from bottom port below cryocooler. Torqued using short arm of allen key (no access for long arm).
    • Borrowed one bolt location for RTD spring clamp mount. Confirmed continuity of coldhead RTD.
  • Repaired pins damaged in removal, and inserted sheaths intended to replace the previous scheme and hopefully prevent pin and socket damage. More detail later :).
  • Installed shields.
    • Applied new Apiezon N thermal grease to bottom flange of inner shield.
    • Re-clamped inner shield to cold plate using a dog clamp array.
    • Placed outer shield in contact with outer shield; rocking percieved when pressing on one quadrant of the annulus.
    • Confirmed continuity of shield RTDs.
  • Installed test mass, still with heater mounted to planar surface as in QIL/2715.
    • Confirmed suspension wires were contacting bare silicon OD surface in middle.
    • Confirmed continuity of heater and of workpiece RTD.
  • Closed up, pumped down, cooled down.
    • Roughing pump on at ~4:20 pm
    • Turbo on with no issues after 15 minute programmed delay
    • Cooling started at 5:15 pm, pressure was a few E-4 torr; no need to wait so long, that was just when we got around to it.

All images are (or soon to be) posted to the QIL Photo Dump.

Attachment 1: D2000310_y-004_section_double_linkage_20220207.png  548 kB  | Hide | Hide all
D2000310_y-004_section_double_linkage_20220207.png
ELOG V3.1.3-